### v1

# From Sufficiency to Reflection: Reinforcement-Guided Thinking Quality in Retrieval-Augmented Reasoning for LLMs

Abstract
Reinforcement learning-based retrieval-augmented generation (RAG) methods enhance the reasoning abilities of large language models (LLMs). However, most rely only on final-answer rewards, overlooking intermediate reasoning quality. This paper analyzes existing RAG reasoning models and identifies three main failure patterns: (1) information insufficiency: failure to retrieve adequate support; (2) faulty reasoning: logical or content-level flaws despite sufficient information; (3) answer–reasoning inconsistency: a valid reasoning chain with a mismatched final answer. We propose TIRESRAG-R1, a novel framework using a think–retrieve–reflect process and a multi-dimensional reward system to improve reasoning and stability. TIRESRAG-R1 introduces: (1) a sufficiency reward to encourage thorough retrieval; (2) a reasoning quality reward to assess rationality and accuracy of the reasoning chain; (3) a reflection reward to detect and revise errors. It also employs a difficulty-aware reweighting strategy and training sample filtering to boost performance on complex tasks. Experiments on four multi-hop QA datasets show TIRESRAG-R1 outperforms prior RAG methods and generalizes well to single-hop tasks [^1] The code and data can be found at https://github.com/probe2/TIRESRAG-R1..
From Sufficiency to Reflection: Reinforcement-Guided Thinking Quality in Retrieval-Augmented Reasoning for LLMs
Jie He1  Víctor Gutiérrez Basulto211footnotemark: 1  and   Jeff Z. Pan1thanks:   Corresponding author 1 School of Informatics, University of Edinburgh, UK 2 School of Computer Science and Informatics, Cardiff University, UK {j.he, j.z.pan}@ed.ac.uk gutierrezbasultov@cardiff.ac.uk
1 Introduction
Large language models (LLMs) Grattafiori et al. (2024); Yang et al. (2025); OpenAI et al. (2024a) have achieved remarkable performance across a wide range of downstream tasks, such as mathematical reasoning Ahn et al. (2024), code generation Jiang et al. (2025), open-domain question answering Kamalloo et al. (2023). A key factor behind this success is chain-of-thought prompting Wei et al. (2022), which enables LLMs to generate intermediate reasoning steps before arriving at a final answer, significantly enhancing performance on reasoning tasks Wang et al. (2023); Pan et al. (2023); Snell et al. (2025). Despite these advances, the internal knowledge of LLMs is not always reliable. For instance, when faced with time-sensitive queries Mousavi et al. (2024) or conflicting evidence within their internal representations Xu et al. (2024), LLMs often produce hallucinations or factual inaccuracies due to outdated or ambiguous information Li et al. (2024a); Wang et al. (2024b).
Figure 1: An example showing different reasoning trajectories for answering a multi-hop query. It compares insufficient information, incorrect predictions, and fully correct reasoning.
The retrieval-augmented generation (RAG) paradigm Gao et al. (2024a) improves factual accuracy and robustness by enabling LLMs to access up-to-date external knowledge. However, standard RAG often treats retrieval and generation as loosely coupled, lacking mechanisms for multi-step reasoning such as query decomposition or identifying knowledge gaps, which limits performance on tasks requiring deeper reasoning.
To overcome this, a growing body of research has turned to reinforcement learning (RL) (a paradigm that has demonstrated strong performance in mathematical reasoning and code generation Shao et al. (2024); Wang et al. (2024a)) to train LLMs for retrieval-augmented reasoning. In this setting, the model learns when and how to retrieve, as well as how to integrate retrieved information into coherent reasoning chains, guided by outcome-level rewards (e.g., answer correctness) Jin et al. (2025a); Song et al. (2025); Sun et al. (2025a). Despite recent progress, outcome-based RL methods still face two significant limitations: (1) they often neglect the quality and validity of intermediate reasoning steps, and (2) they lack fine-grained feedback to guide the reasoning process during multi-step tasks. As a result, models may learn incorrect reasoning paths that degrade answer quality, or they may arrive at correct answers through flawed reasoning, thereby compromising interpretability.
In this paper, to address these challenges, we begin by evaluating the performance of RAG-based reasoning models trained with outcome-based rewards. Through human analysis on examples from multi-hop reasoning datasets (e.g., MuSiQue Trivedi et al. (2022)), as shown in Fig. 1, we identify three primary bottlenecks limiting model accuracy: (1) the model retrieves sufficient information but still fails to produce the correct answer; (2) reasoning is prematurely interrupted, resulting in inadequate retrieval; and (3) the reasoning chain is correct, yet the final answer is incorrect. Based on these insights, we propose TIRESRAG-R1, an approach that enhances reasoning chain quality by incorporating rewards focused on sufficiency and reasoning coherence. Additionally, we introduce a thinking–retrieval–reflection framework, motivated by the principle that models should first ensure the integrity of their reasoning chains. If the final answer is incorrect despite a valid chain, the model engages in reflection to revise its response.
TIRESRAG-R1 enhances RAG-based reasoning by introducing a thinking–retrieval–reflection framework, where the model can decide post-answer whether to reflect and revise its output through an additional round of reasoning and retrieval. To guide learning, TIRESRAG-R1 assigns multi-dimensional reward signals to each reasoning trajectory, including: an answer reward, a sufficiency reward to encourage retrieval of adequate evidence, a reasoning quality reward evaluating logical coherence, alignment, error awareness, and conciseness, and a reflection reward that promotes correction of wrong answers while discouraging unnecessary changes. To address the varying difficulty of retrieval tasks, we propose a difficulty-aware advantage reweighting strategy that adjusts reward weights based on the ease of retrieving sufficient evidence, and an advantage filtering mechanism that excludes trivial examples (where all responses are correct or incorrect) to stabilize RL training. Together, these components enable more robust and interpretable reasoning, significantly outperforming traditional outcome-based methods across multiple QA benchmarks.
Our contributions can be summarized as follows:
  • We are the first to define overthinking and underthinking in RAG and conduct comprehensive analysis to reveal the main bottlenecks faced by current RL-trained RAG reasoning models.
  • We propose TIRESRAG-R1, which enhances reasoning by introducing a reflection mechanism and fine-grained multi-dimensional rewards to improve both reasoning chains and answer accuracy. We also propose a difficulty-aware advantage reweighting strategy with no additional computational cost to guide the model toward optimizing on harder questions.
  • Our experiments show that TIRESRAG-R1 outperforms multiple state-of-the-art RAG methods on several open-ended QA datasets, with an average accuracy improvement of 5.8%. Further ablation studies and analysis confirm the effectiveness of our fine-grained rewards.
2 Related Work
Retrieval-Augmented Generation. RAG has been widely adopted to mitigate hallucination, domain knowledge gaps, and temporal staleness Ayala and Bechard (2024); Siriwardhana et al. (2023); Gade and Jetcheva (2024). Traditional RAG typically follows a static retrieve-then-generate pipeline Lewis et al. (2020); Guu et al. (2020), which is effective for open-domain QA but often struggles with multi-hop reasoning, latent constraints, and ambiguous queries Tang and Yang (2024); Li et al. (2024b); Chan et al. (2024). Recent work has proposed more cognitively inspired architectures, such as AdaptiveRAG Jeong et al. (2024) using query classification to choose retrieval strategies. In parallel, modular and hybrid frameworks Gao et al. (2024b); Zhang et al. (2024); Zhou et al. (2024) integrate evidence aggregation, verification, and query rewriting, reflecting a deeper interaction between retrieval and reasoning.
Reinforcement Learning for LLM Reasoning. Reinforcement learning (RL) has recently been explored as a way to improve multi-step reasoning in LLMs. Models such as GPT-o1 and DeepSeek-R1 OpenAI et al. (2024b); DeepSeek-AI et al. (2025) demonstrate RL-based training for structured reasoning. Follow-up work (e.g., SimpleRL-Zoo Zeng et al. (2025), Open-Reasoner-Zero Hu et al. (2025c), and Light-R1 Wen et al. (2025)) further leverage curriculum design, reward shaping, and improved optimization algorithms such as Dr GRPO Liu et al. (2025) to enhance verifiability and coherence.
Reinforcement Learning for RAG Reasoning. Recent studies combine RL with RAG to dynamically guide when and what to retrieve Song et al. (2025); Jin et al. (2025a); Chen et al. (2025). Early approaches rely primarily on outcome-level rewards, while newer ones introduce process-level rewards Wang et al. (2025); Sha et al. (2025); Zhang et al. (2025b). For instance, R3-RAG Li et al. (2025b) computes document relevance at each step to refine search strategies, and others enrich the “search–think–answer” pipeline with additional fields to prompt deeper reflection Ren et al. (2025); Shi et al. (2025). In contrast, our work directly rewards the model’s reasoning process by measuring the sufficiency of retrieved documents, enabling more adaptive retrieval-reasoning coordination.
For more details on related work, please see Appendix A.
3 Prelimary Study
Although R1-like RAG models demonstrate stronger reasoning capabilities than traditional RAG models in QA tasks, questions remain about the faithfulness of their reasoning processes, both in relation to the provided context and in how well their final answers reflect that reasoning. This section focuses on the following central question: In current reasoning-oriented RAG models tackling multi-hop tasks, what is the correlation between the predicted answers, the retrieved documents, and the generated reasoning text?
Category Datasets 2Wiki MuSiQue
Corr. Incorr. Corr. Incorr.
Overthinking 13 25 13 31
Good thinking 232 89 80 96
Underthinking 16 125 0 280
Table 1: Distribution of correct (Corr.) and incorrect (Incorr.) predictions across different reasoning categories on 2Wiki and MuSiQue datasets.
3.1 Preliminary Analysis
We trained a reasoning-capable RAG model using the GRPO algorithm from DeepSeek-R1, following the data setup of R1-Searcher. We used Qwen-2.5-3B-Instruct with a local retriever and evaluated on in-domain 2Wikimultihopqa (2Wiki) and out-of-domain Musique test sets.
In what follows, we will denote an example as $E=(Q,RD,A)$, where $Q$ is the query, $RD$ the combined reasoning steps and retrieved documents, and $A$ the predicted or gold answer. Regarding reasoning behavior, we categorize the model’s reasoning into three cases: overthinking (gold answer could be inferred before the final step), good thinking (gold answer emerges exactly at the final step), and underthinking (model stops without sufficient content). Following prior work Joren et al. (2025), we use an LLM to assess whether $Q$ and $RD$ are sufficient to derive the gold answer, but unlike Joren et al. (2025), we also provide the gold answer to reduce misjudgments. Details of fine-tuning setup and experimental analysis settings are all provided in Appendix J.
Context Sufficiency Error type Case Ratio
Insufficient Knowledge retrieved incorrect Question: What is the place of birth of the founder of Leister Technologies? Model output:<search> founder of Leister Technologies </search><think> Louis Leitz (founder of Esselte Leitz) … therefore answer Ingersheim </think> <answer> Ingersheim </answer> Gold Answer: Wilhelmshaven 30.2%
Insufficient Thinking incorrect Question: Where was the place of burial of John De Ferrers, 1st Baron Ferrers of Chartley’s father? Model output:<search> John De Ferrers, 1st Baron Ferrers of Chartley father’s burial place </search><think> Linked to Sir Humphrey Ferrers of Tamworth Castle and concluded burial at Tamworth Church </think> <answer> Tamworth Church </answer> Gold Answer: Staffordshire 22.1%
Sufficient Thinking process mismatch evidence Question: Who is Gaius Rubellius Blandus’s father-in-law? Model output:<search> Gaius Rubellius Blandus father-in-law </search><think> Married Julia (granddaughter of Tiberius) $\Rightarrow$ therefore father-in-law is Tiberius </think> <answer> Tiberius </answer> Gold Answer: Drusus Julius Caesar 16.4%
Sufficient Thinking correct but incorrect answer Question: Which film has the director died earlier, Girls School Screamers or Lucky Numbers? Model output:<search> director of Girls School Screamers </search><search> director of Lucky Numbers </search><think> Found John P. Finnegan (died 2007) and Nora Ephron (died 2012) $\Rightarrow$ therefore the director of Lucky Numbers director died earlier </think> <answer>Girls School Screamers</answer> Gold Answer: Lucky Numbers 5.7%
Sufficient Thinking process match evidence but thinking process is incorrect Question: Which film has the director who was born first, Nasib Si Labu Labi or Solo Per Te Lucia? Model output:<search> director of Nasib Si Labu Labi birth year </search><search> director of Solo Per Te Lucia birth year </search><think> P. Ramlee (born 1936) vs Franco Rossi (born 1923) $\Rightarrow$ therefore Nasib Si Labu Labi director was born first </think><answer> Nasib Si Labu Labi </answer> Gold Answer: Solo Per Te Lucia 25.6%
Table 2: Fine-grained error analysis with examples.
We report sufficiency judgment results separately based on whether the model’s prediction is correct. The results are shown in Table 1. We can observe that among correctly predicted examples, overthinking and underthinking account for very small proportions, while good thinking accounts for 86% in Musique and 89% in 2Wiki. This indicates that for correctly predicted examples, the model’s reasoning process is generally just right and sufficient. Among incorrectly predicted examples, underthinking accounts for a large proportion. This is unsurprising, as the model did not retrieve relevant documents and thus lacked enough information to answer the question, so answering incorrectly is expected. Surprisingly, overthinking and good thinking together account for 48% in Musique and 31% in 2Wiki. This shows that the reasoning process was sufficient, but the model still failed to produce the correct answer.
3.2 Fine-grained Analysis
Since overthinking accounts for a small proportion among predictions, we further analyzed the incorrectly predicted examples belonging to good thinking and overthinking to determine specific error causes. We manually reviewed these examples and classified them into five error types, as shown in Table 2.
We observe the following. When the reasoning process was insufficient to fully answer the question, we found two error types: (i) the model mistakenly believes the retrieved content matches the question entity and therefore stops further retrieval and outputs an answer, while in fact the retrieved content is irrelevant; (ii) the model only obtained partial content relevant to the question and ended the reasoning, which is a reasoning omission. When the reasoning process was sufficient to answer the question, a prominent problem was that the model failed to follow evidence, indicating difficulty in understanding retrieved content and integrating it into reasoning Shi et al. (2025); Li et al. (2025a). Another issue was that the reasoning process was correct but the final answer was wrong, or the model showed correct evidence but failed to reason correctly from it.
Based on these findings, we categorize the failure cases of RAG reasoning on part of the test sets into three major types: retrieval failure, reasoning failure, and answer failure. Motivated by this observation, we propose TIRESRAG-R1, which aims to enhance the model’s awareness of retrieval sufficiency while improving both its reasoning chain and final answer quality.
Figure 2: Overall architecture of TIRESRAG-R1, which integrates search, reasoning, and reflection with a dynamic reward design to guide reinforcement learning.
4 Method
In this section, we first introduce our GRPO-based thinking–search–answer–reflect pipeline. Then, we elaborate on the reward design, which is the core of our RL algorithm. On top of the standard answer reward, we carefully add three additional reward signals: thinking reward, sufficient reward, and reflect reward. Finally, we observe that during training there exists a class of “extreme” examples that are either answered correctly by the model in all rollouts or answered incorrectly in all rollouts. Such examples contribute very limited training signals and may even introduce noise, affecting training stability. To address this issue, we introduce two optimization mechanisms: a difficulty-aware advantage reweighting strategy and a group filtering mechanism that filters out those extreme problems. A complete algorithmic description can be found in Appendix K.
4.1 Trajectory Generation with Search and Reflect
Given a question $q$, we first prompt the LLM $\pi_{\theta}$ to generate a long chain of thought. During this thinking process, the model is encouraged to trigger search operations in order to use external document information. When the model decides that a search is needed, it terminates the current thinking step. The reasoning text so far is wrapped with <think> </think> tags, and the search query is wrapped with <search> </search> tags. We extract the search query and feed it into the retriever $\pi_{\text{ret}}$ to obtain $k$ relevant documents $D_{k}=\{d_{1},\dots,d_{k}\}$. These documents are wrapped with <information> </information> tags and appended back to the trajectory. The model then continues reasoning based on the retrieved documents. This process of thinking and searching repeats until the model judges that enough information has been collected to generate the final answer. At that point, the answer is wrapped with <answer> </answer> tags. However, according to our prompt design, the model is instructed to reflect on the generated answer, potentially entering another cycle of thinking and searching before producing a second answer wrapped again with <answer> </answer>. This reflection mechanism is specifically introduced to address the issue discussed in Section  3, namely that the reasoning process may be correct but the first generated answer is wrong. In the final trajectory, the content inside the last <answer> </answer> is taken as the model’s predicted answer.
4.2 GRPO for Training
As shown in Figure 2, we adopt the group relative policy optimization (GRPO) algorithm DeepSeek-AI et al. (2025) for RL training. For each query in GRPO, a group of $G$ rollout trajectories, as described in Section 4.1, is generated using the current policy $\pi_{\theta}^{\text{old}}$. Here $\pi_{\theta}^{\text{old}}$ also serves as a frozen reference model initialized with the same parameters as the policy model. The GRPO algorithm uses the following optimization objective to update the policy:
$J_{\text{GRPO}}(\theta)={}$ $\mathbb{E}_{q\sim Q,\{o_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\min$
$\Bigg{(}\frac{\pi_{\theta}(o_{i,t}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q)}A_{i},$ $\text{clip}\!\left(\frac{\pi_{\theta}(o_{i,t}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q)},1-\epsilon,1+\epsilon\right)A_{i}\Bigg{)}\Bigg{]}$
where $\epsilon$ is the clipping hyper-parameter and $\beta$ is the KL-divergence penalty coefficient. The advantage $A_{i}$ for each response is computed as $A_{i}^{=}\tfrac{r_{i}-\text{mean}(\{r_{j}\}_{j=1}^{G})}{\text{std}(\{r_{j}\}_{j=1}^{G})}$, where $\{r_{i}\}_{i=1}^{G}$ are the rewards from the group. The detailed definition of each reward is introduced in Section 4.3.
4.3 Reward Design
As described in Section 3, we aim to achieve three major goals: (1) encourage the policy model to search sufficiently before generating an answer, (2) improve the quality of the reasoning chain and make it more consistent with the retrieved information, and (3) enhance the model’s ability to correct answers through reflection. To achieve these goals, we adopt a mixed reward function with the following components:
Answer reward measures the match between the predicted answer and the gold answer. Following prior work Jin et al. (2025a), we use the F1 score as the answer reward to avoid reward hacking issues observed with exact match (EM):
$R^{A}=F_{1}(a,a^{*}).$ (1)
Sufficient reward measures whether the reasoning trajectory $RD$ provides enough information to support the gold answer. Following Section 3, we use a locally deployed LLM to score $(Q,RD,\text{gold answer})$ with 0/1, where 1 means sufficient and 0 means insufficient:
$R^{S}=\begin{cases}1.0,&\text{if $(Q,RD,o_{g})$ is sufficient},\\ 0.0,&\text{otherwise}.\end{cases}$ (2)
Thinking reward measures the quality of the thinking part of the trajectory. The locally deployed LLM is prompted to evaluate logical soundness, alignment with retrieved content, error awareness, and concise accuracy. It outputs a score in [0,1][0,1][ 0 , 1 ]:
$R^{T}\in[0,1].$ (3)
Reflect reward encourages the model to revise wrong answers into correct ones. We extract each intermediate answer $a_{1},a_{2},\dots$ from the trajectory, compare each with the gold answer using an accuracy score, and assign rewards as follows: a positive reward when reflection corrects an incorrect answer, a negative reward when it replaces a correct answer with an incorrect one, and zero otherwise.
$R^{R}=\begin{cases}+1.0,&\text{if }CEM(a_{1},a^{*})=0\text{ and }CEM(a_{2},a^{*})=1,\\ -1.0,&\text{if }CEM(a_{1},a^{*})=1\text{ and }CEM(a_{2},a^{*})=0,\\ 0.0,&\text{otherwise}.\end{cases}$ (4)
Dynamic Weight of Reward. Because of the mixed reward mechanism, the model could overfit to auxiliary signals such as sufficient reward or thinking reward and ignore answer accuracy. We design a dynamic weight schedule $a_{t}=\tfrac{1}{1+\exp\!\big{(}\tfrac{t-0.9T}{10}\big{)}}$, where $t$ denotes the current training step and $T$ denotes the total number of training steps. This schedule gradually shifts focus from auxiliary reasoning rewards to answer accuracy as training proceeds.
Finally, the overall reward is computed as
$R^{\text{sum}}=R^{A}+a_{t}\bigl{(}w_{t}R^{T}+w_{s}R^{S}+w_{r}R^{R}\bigr{)},$ (5)
4.4 Optimization Strategy
Difficulty-aware Resampling. Despite employing a mixed reward mechanism, we observed that different rollout groups can yield very different raw rewards but almost identical normalized advantages (e.g., [0.8,0.85,0.9,0.95,1.0][0.8,0.85,0.9,0.95,1.0][ 0.8 , 0.85 , 0.9 , 0.95 , 1.0 ] vs. [0.0,0.05,0.1,0.15,0.2][0.0,0.05,0.1,0.15,0.2][ 0.0 , 0.05 , 0.1 , 0.15 , 0.2 ]). This phenomenon causes simple problems to receive the same optimization emphasis as hard problems. Following Zhang and Zuo (2025), we introduce a difficulty-aware advantage reweighting strategy. For each sample, we estimate the problem difficulty by computing the average sufficient-reward score across all its rollouts, denoted as $R_{S}^{\text{avg}}$. The weight function is defined as
$W(R^{S}_{\text{avg}})=A+\frac{B-A}{1+\exp\!\big{[}k(R^{S}_{\text{avg}}-\rho_{0})\big{]}},$ (6)
where $A,B,\rho_{0},k$ are tunable hyper-parameters controlling the sensitivity of the reweighting. $R^{s}_{\text{avg}}$ is the average sufficient-reward score of all rollouts associated with question $q$.
Consistency Penalty. Because we use a mixed reward mechanism, it is possible for a trajectory to receive a high reasoning reward but a low answer reward. To encourage consistent trajectories where high reasoning quality aligns with high answer accuracy, we add a small penalty term to discourage such inconsistencies:
$A_{i}^{P}=-\lambda_{p}A_{i}^{T}\cdot A_{i}^{T}\cdot A_{i}^{A},\quad\text{if }A_{i}^{S}\cdot A_{i}^{T}\cdot A_{i}^{A}<0.$ (7)
where $A_{i}^{S}$, $A_{i}^{T}$, and $A_{i}^{A}$ denote the group-normalized values of the sufficient reward, the thinking reward, and the answer reward, respectively. The final difficulty-aware advantage is then expressed inline as $A_{i}^{\prime}=(A_{i}-A_{i}^{P})\cdot W(R^{s}_{\text{avg}})$, where $A_{i}$ is the original normalized advantage for rollout $i$.
Group Filtering. As noted in Jin et al. (2025a), applying GRPO to RAG tasks can lead to late-stage collapse, where training rewards drop to zero. We also observe this phenomenon: in later stages, many queries have groups where all rollouts are either completely correct or completely wrong, leading to zero advantage and noisy updates. To mitigate this, we introduce a simple filtering mechanism that removes these saturated queries from the training set.
Method Hotpotqa 2wikimultihopqa Musique Bamboogle
EM F1 Judge EM F1 Judge EM F1 Judge EM F1 Judge
Direct Generation 18.0 23.5 24.2 19.0 23.6 23.0 3.6 9.0 6.0 20.8 30.7 28.8
COT 18.2 24.4 25.2 21.0 25.7 25.0 4.4 10.9 8.8 21.6 31.3 28.8
Naive RAG 29.4 40.7 43.4 26.0 30.5 29.8 5.2 11.0 8.0 18.4 27.0 22.4
Sure 20.2 27.2 32.0 21.0 24.6 26.2 3.0 8.5 5.2 10.4 17.9 12.8
IRCOT 22.2 32.7 38.6 17.4 23.4 26.0 5.2 10.8 8.4 13.6 23.8 24.0
Self-ask 14.4 24.6 39.0 14.8 19.6 25.0 4.0 10.3 7.4 9.6 22.3 18.4
RAG with Agentic Search 7.0 10.1 10.4 10.4 12.3 12.4 1.4 6.8 3.4 9.6 13.3 12.0
Search-o1 12.4 17.6 17.2 17.0 19.6 18.8 3.4 8.8 6.4 14.4 23.0 18.4
SFT 15.8 20.1 18.2 28.4 31.1 30.0 2.2 9.6 3.8 8.0 15.8 8.8
SimpleDeepSearcher 34.4 45.3 47.6 39.6 46.5 47.8 12.4 20.5 18.8 33.6 44.1 42.4
ReSearch-Base 28.8 38.4 36.2 37.2 40.7 40.0 14.4 24.2 17.4 34.8 45.5 37.2
ReSearch-Instruct 30.8 41.7 43.6 38.0 42.0 41.6 14.2 23.8 18.0 34.8 47.1 42.0
R1-search-Base 30.8 40.7 48.0 37.2 41.1 40.2 13.6 23.9 17.0 33.2 41.3 37.2
R1-search-Instruct 31.2 42.2 43.4 42.6 47.1 46.2 15.4 25.3 18.6 33.2 43.5 39.6
Search-R1-Base 35.4 50.1 51.2 44.0 50.9 51.4 14.8 26.0 22.6 38.4 50.9 48.4
Search-R1-Instruct 37.4 49.3 50.4 47.6 51.7 53.4 16.2 23.7 21.0 40.2 50.3 47.4
LeTS-Instruct* 37.1 - 55.2 41.0 - 47.5 17.5 - 26.9 38.4 - 51.2
TIRESRAG-R1-Base 41.0 53.0 56.4 52.4 58.4 60.2 16.2 26.9 24.4 40.4 52.6 50.0
TIRESRAG-R1-Instruct 41.0 54.2 56.0 52.8 59.6 61.4 19.4 30.0 27.4 44.0 54.7 52.8
Table 3: Main experimental results on HotpotQA, 2WikiMultiHopQA, Musique, and Bamboogle.
5 Experiments
Experimental Setup. Following the evaluation protocol of R1-Searcher, we assess performance on four widely used and challenging multi-hop QA datasets: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), Bamboogle Press et al. (2023), and Musique Trivedi et al. (2022). For HotpotQA, 2WikiMultiHopQA, and Musique, we adopt the test splits released by R1-Searcher, each containing 500 examples. For Bamboogle, we use the full set of 125 test examples. For training, we use the dataset provided by R1-Searcher, which includes 4,561 examples from the HotpotQA training set and 3,587 examples from the 2WikiMultiHopQA training set.
Evaluation Metrics. We evaluate model performance using four metrics: Exact Match (EM), F1, LLM-as-Judge, and Cover Exact Match (CEM). EM checks exact matches, F1 accounts for partial overlaps, LLM-as-Judge (via GPT-4o) assesses semantic correctness, and CEM measures whether the gold answer is covered. For more detailed settings, please refer to App. C.2.
Baselines. We compare TIRESRAG-R1 with 14 representative baselines spanning four categories: (1) Naive prompt methods: Direct, COT, and R1-based; (2) Retrieval-augmented prompt methods: Naive RAG, Agentic-R1, Search-o1, SURE, IRCOT, Self-Ask, and RQRAG; (3) SFT methods: SFT and SimpleDeepSearcher; (4) RL methods: Search-R1, R1-Searcher, Research, and the process-reward method LeTS. For detailed descriptions and configurations of these baselines, please refer to the App. C.3.
5.1 Main Results
Table 3 shows that TIRESRAG-R1 achieves superior or competitive performance on all four complex multi-hop datasets when trained on either Qwen-2.5-3B-Base or Qwen-2.5-3B-Instruct. Our key findings are as follows:
(1) For the small 3B models, prompt-based methods generally perform poorly. For instance, Search-o1 even underperforms naive RAG, as 3B models struggle to interpret instructions and generate effective retrieval queries without fine-tuning.
(2) Incorporating format rewards may not be optimal. Compared to Search-R1 (answer reward only), Research and R1-Searcher exhibit average EM drops of 5.13% and 4.60%, respectively. We attribute this decline to format-based learning, which reduces the exploration capacity of the 3B models and weakens useful reward signals when answers are correct but deviate slightly in format.
(3) TIRESRAG-R1 delivers substantial performance gains. On in-domain datasets HotpotQA and 2WikiMultiHopQA, it outperforms Search-R1 by average EM margins of 4.7% and 7.0%, respectively. For out-of-domain datasets Musique and Bamboogle, we observe improvements of 5.3% and 4.6%. Compared to LeTS, TIRESRAG-R1 achieves additional gains of 5.8 and 4.2 points in EM and LLM-as-Judge, respectively. These results indicate that our approach effectively learns higher-quality reasoning chains and generalizes well to unseen domains.
Method Hotpotqa 2wikimultihopqa Musique Bamboogle
EM F1 CEM EM F1 CEM EM F1 CEM EM F1 CEM
Ours-Base+GRPO 41.0 53.0 47.8 52.4 58.4 59.0 16.2 26.9 21.4 40.4 52.6 43.6
Ours-Instruct+GRPO 41.0 54.2 46.0 52.8 59.6 60.8 19.4 30.0 23.2 44.0 54.8 47.2
Ours-Base+Reinforce++ 42.2 55.9 49.0 52.0 58.6 61.2 16.6 28.7 22.2 36.8 50.7 44.0
Ours-Instruct+Reinforce++ 39.6 53.9 49.8 46.2 55.0 58.4 15.4 25.4 20.6 37.6 48.8 42.4
Table 4: Comparison between GRPO and Reinforce++ across datasets.
Datasets HotpotQA 2WikiMultiHopQA MusiQue Bamboogle Average
Thinking Length
Naive grpo 318.7 293.1 467.7 215.2 323.7
Search-R1-Instruct 258.4 331.3 360.8 194.3 286.2
TIRESRAG-R1-Instruct 252.1 312.1 303.6 229.2 274.3
Search Steps
Naive grpo 2.7 2.9 3.9 2.2 2.93
Search-R1-Instruct 2.3 2.3 2.8 2.2 2.40
TIRESRAG-R1-Instruct 2.1 2.7 2.6 2.0 2.35
Table 5: Comparison of average thinking length and search steps across datasets.
Method NQ PopQA TriviaQA
COT 10.5 9.7 7.8
Sure 25.5 30.4 13.4
SimpleDeepSearcher 33.4 38.9 59.6
ReSearch-Instruct 35.8 41.8 58.4
TIRESRAG-R1-Instruct 38.0 43.0 60.0
Table 6: Generalization results on single-hop benchmarks.
5.2 Analysis
We perform a comprehensive analysis to better understand the factors influencing our method’s effectiveness. Below, we highlight the most significant findings; for a more detailed discussion and a case study, please refer to Appendix G.
Different RL Methods. To assess the generality of our training strategy, we also apply it to Reinforce++ Hu et al. (2025a) that does not use a critic model. The key difference between Reinforce++ and GRPO is that it normalizes advantages across the entire batch rather than within group rollouts. We train both Qwen2.5-3B-Base and Instruct models under identical settings. As shown in Table 4, compared with GRPO’s group-normalization strategy, Reinforce++ performs better on in-domain datasets (HotpotQA, 2WikiMultiHopQA) but worse on out-of-domain datasets (Musique, Bamboogle), which is consistent with the findings in R1-Searcher.
Efficiency Analysis. We compare search counts and thought lengths with naive GRPO and Search-R1. As shown in Table 5, despite encouraging reflection and sufficient information gathering, our method does not hurt efficiency. Compared to naive GRPO, our method reduces average search count by 0.58 points and token generation by 49.4 points. Compared to Search-R1, search count decreases by 0.05 points and token generation by 12 points. This indicates that RL-trained Agent-RAG models can achieve improved reasoning quality without sacrificing efficiency.
5.3 Generalization on Single-Hop Benchmarks
Since our primary training data consist of multi-hop QA examples, we also evaluate on single-hop tasks to assess generalization. We use three widely adopted single-hop QA benchmarks: NQ, PopQA, and TriviaQA and present comparisons with the best baseline in each category. More comprehensive results can be found in the App. H. Table 6 shows that our method consistently outperforms all baselines across all metrics. On NQ, it achieves EM score of 38.0, exceeding the strongest baseline (ReSearch-Instruct) by 2.2 points. On PopQA, we obtain 43.0, surpassing baselines by 1.5 points. For TriviaQA, we achieve 60.0, with an improvement of 2.1 points. These results demonstrate that, despite being trained primarily on multi-hop data, our method generalizes effectively to single-hop tasks, highlighting its robustness across reasoning types and task distribution.
6 Conclusion
In this work, we reveal the limitations of current outcome-supervised RL-trained RAG reasoning models. To address these issues, we propose TIRESRAG-R1, which uses a novel think–search–reflect paradigm, explicitly encouraging models to reflect on uncertain answers. We design multiple reward functions to improve document retrieval sufficiency and reasoning quality, and introduce a difficulty-aware advantage strategy to provide stronger learning signals on hard problems. Comprehensive evaluations show that TIRESRAG-R1 outperforms existing methods on four multi-hop QA benchmarks. Further analysis highlights that our method is compatible with different RL algorithms and exhibits strong generalization.
Limitations
Model scaling. Our experiments are conducted only on Qwen2.5-3B due to computational constraints. Although TIRESRAG-R1 shows promising results on smaller models, its effectiveness on larger architectures (e.g., Qwen2.5-7B) remains unexplored.
Reward modeling. We use Qwen3-8B for sufficient and thinking scoring. Although it provides accurate signals, using a stronger model such as GPT-4o, or a specialized reward model fine-tuned on domain-specific data, may further improve performance. Exploring more accurate reward models (e.g., as suggested in recent thinking-reward literature) is a promising direction.
Reflection signal sparsity. While our reflection mechanism corrects some wrong answers, the number of training examples requiring reflection is limited, reducing useful learning signals. Future work could synthesize reflection-rich data for SFT before applying TIRESRAG-R1 to further improve the model’s ability to reflect effectively.
Appendix A Related Work
Retrieval-Augmented Generation. RAG has emerged as a prominent framework to augment LLMs with external knowledge, aiming to mitigate issues such as hallucination Ayala and Bechard (2024), domain incompleteness Siriwardhana et al. (2023), and temporal staleness Gade and Jetcheva (2024). Conventional RAG systems adopt a static retrieve-then-generate paradigm, where a retriever first fetches top-ranked documents given an input query, and a generator conditions its output on the retrieved context Lewis et al. (2020); Guu et al. (2020). While this structure has proven effective for factual QA and open-domain generation, it often falls short when confronted with complex reasoning tasks involving multi-hop dependencies Tang and Yang (2024), latent constraints Li et al. (2024b), or ambiguous query intents Chan et al. (2024). To overcome these limitations, recent works have proposed more cognitively informed RAG architectures. For example, AdaptiveRAG Jeong et al. (2024) uses query classification to trigger different retrieval strategies, PlanRAG Lee et al. (2024) decomposes tasks into executable plans for targeted retrieval, and ITER-RETGEN Shao et al. (2023) incorporates intermediate generation to iteratively reformulate queries. In parallel, modular and hybrid RAG frameworks Gao et al. (2024b); Zhang et al. (2024); Zhou et al. (2024) have introduced componentized systems that integrate query rewriting, evidence aggregation, and verification in sequential or recursive pipelines. These advances suggest that effective RAG increasingly requires not just better retrieval quality, but dynamic, context-aware reasoning to inform when, what, and how to retrieve. This growing interdependence between retrieval and reasoning sets the stage for more adaptive mechanisms (particularly those that go beyond pre-defined rules) highlighting the need for learning-based control in complex RAG workflows.
Reinforcement Learning for LLM Reasoning. Driven by the growing need for LLMs to perform complex and reliable reasoning across diverse tasks, recent research has turned to reinforcement learning as a promising paradigm to enhance their reasoning capabilities. The release of GPT-o1 OpenAI et al. (2024b) and DeepSeek-R1 DeepSeek-AI et al. (2025) marked a shift toward training LLMs that exhibit structured, multi-step reasoning through RL-based objectives. Early efforts such as SimpleRL-Zoo Zeng et al. (2025) and Open-Reasoner-Zero Hu et al. (2025c) explored direct RL fine-tuning from base models, eliminating the reliance on extensive supervised instruction tuning. Building on this foundation, approaches like DeepScaler Meng et al. (2024) and Light-R1 Wen et al. (2025) introduced cold-start datasets and reward schemes explicitly designed to promote step-by-step thinking and verifiable inferences. In parallel, improvements to RL algorithms, such as Dr GRPO Liu et al. (2025), refined policy optimization to better align with the cognitive demands of long-form reasoning.
Reinforcement Learning for RAG Reasoning. There is a growing body of work focused on bringing RL-based reasoning into the retrieval-augmented generation framework Song et al. (2025); Jin et al. (2025a); Chen et al. (2025). Inspired by DeepSeek-R1, these approaches use regularized rewards to encourage the model to think and generate retrieval queries to search external corpora. However, these methods only consider the final outcome reward signal to train the model, which is relatively simple, and do not deeply optimize according to the characteristics of the RAG task itself. In view of this, many works incorporate the model’s thinking process into the reward calculation Wang et al. (2025); Sha et al. (2025); Zhang et al. (2025b). For example, Zhao et al. (2025) uses different models to generate answers based on the same evidence in order to calculate evidence quality, while also alleviating answer bias caused by the preference of the policy model itself. R3-RAG Li et al. (2025b) calculates the relevance between each retrieved document and the question at every step, attempting to improve the model’s search strategy through fine-grained process rewards. Different from these works, our work directly measures the sufficiency between all retrieved documents and the question, in order to enhance the model’s awareness of searching less or more. In addition, some works attempt to add new information fields to the “search–thinking–answer” pipeline proposed by search-r1-type methods to prompt the model to think more about the documents Ren et al. (2025). For example, Shi et al. (2025) lets the model refine the retrieved documents during the reasoning process before proceeding to thinking. Our work, unlike in these methods, directly rewards the model’s thinking process, allowing the model to optimize its reasoning process, rather than manually adding thinking rules.
Appendix B Experimental Setup for Preliminary Studies
We trained a reasoning-capable RAG model using the GRPO algorithm from Deepseek-R1. Following R1-Searcher, we selected part of the training data from HotpotQA Yang et al. (2018) and 2WikiMultiHopQA Ho et al. (2020), with a total of 8,148 examples. We used the Qwen-2.5-3B-Instruct model for training. For the retriever, we deployed a system locally based on the BGE-large-en-v1.5 retrieval, with the retrieval corpus from English Wikipedia provided by KILT Petroni et al. (2021) in 2019. The training prompt can be found in Appendix J. We trained the model for one epoch (detailed experimental settings are the same as those in Section 5.) and evaluated the model on the in-domain 2Wiki test set and the out-of-domain Musique test set. After obtaining predictions on 500 test examples from each dataset, we used GPT-4o to evaluate the correlation between the model’s predictions and the retrieved content plus the reasoning process.
We introduce notation for a generic open-domain question–reasoning setting, assuming sufficient contextual information. Consider an example represented as $E=(Q,RD,A)$, where $Q$ is the query, $RD$ is the combined reasoning process and retrieved documents, and $A$ is either the model’s predicted answer or the gold answer. We define $RD=\{R_{1},D_{1},\dots,R_{i},D_{i},R_{i+1}\}_{i=1}^{n}$, where $n$ is the number of sub-questions generated by the model during reasoning, $R_{i}$ denotes the intermediate reasoning step prior to document retrieval, and $D_{i}$ represents the set of documents retrieved in response to the sub-question formulated in $R_{i}$.
We define three cases: (1) Overthinking: the model produces too many reasoning steps, meaning that the gold answer could already be inferred at some step $R_{i}$ with $i<n$. (2) Good thinking: the model obtains sufficient content exactly before the final reasoning step, i.e., the gold answer can only be inferred at step $R_{n+1}$. (3) Underthinking: the model stops reasoning without obtaining sufficient content in $RD$, and still outputs an answer. Since Joren et al. (2025) has shown enabling LLMs to judge whether the provided context is adequate for answering a question is effective, we directly input $Q$, $RD$, and the gold answer into an LLM to assess whether $Q$ and $RD$ together are sufficient to derive the correct answer (see prompt in App.J). Importantly, we include the gold answer in the input, unlike what Joren et al. (2025) does. which uses only $Q$ and $RD$. This is because we found that omitting the answer leads to more errors: when $RD$ is actually insufficient, the model is more likely to incorrectly judge it as sufficient.
Appendix C Implementation Details and Baselines
C.1 Implementation Details.
[…]
C.2 Evaluation Metrics
For open-ended evaluation, we report three widely used metrics: Exact Match (EM), F1 score (aligned with the RL training answer reward), and LLM-as-Judge. EM measures whether the ground-truth answer exactly matches the model prediction. F1 is computed between predicted and gold answers to handle partial overlap. For LLM-as-Judge, we use GPT-4o to evaluate the semantic correctness of the prediction given the question and supporting evidence. The prompts used for LLM-as-Judge are listed in App. J. Due to the high cost of using LLM-as-a-Judge, we only adopt this metric in the main experiments. For the other experiments, we use Cover Exact Match (CEM), which assesses whether the ground-truth answer is included in the predicted answer.
C.3 Baselines.
To evaluate the effectiveness of our proposed TIRESRAG-R1 method, we implement and compare against 14 baselines, grouped into four categories:
(1) Naive prompt methods: Direct answers questions directly using its own parametric knowledge. COT is instructed to produce a chain of thought before the final answer. R1-based uses a distilled DeepSeek-Qwen-3B reasoning model to first reason then answer. We implement all three naive prompt methods ourselves.
(2) Retrieval-augmented prompt methods: Naive RAG extends Direct by retrieving documents for the query and appending them as additional input before direct answering. Agentic-R1 Li et al. (2025a) enables the model to autonomously retrieve external knowledge when needed while avoiding interference from irrelevant content. Search-o1 Li et al. (2025a) introduces a Reason-in-Documents module that condenses retrieved content into coherent reasoning steps, iteratively guiding the model to the final answer. SURE Kim et al. (2024) generates and evaluates summaries of retrieved passages for multiple answer candidates. IRCOT Trivedi et al. (2023) interleaves retrieval with Chain-of-Thought reasoning. Self-Ask Press et al. (2023) improves multi-hop reasoning by decomposing the original question into intermediate sub-questions that are answered before final prediction. RQRAG Chan et al. (2024) learns to explicitly refine queries through rewriting, decomposition, and disambiguation. For these retrieval-augmented methods, we use the implementations provided in flashrag.
(3) SFT methods: SFT fine-tunes the model directly on training pairs of questions and gold answers. During training, we only input the question and instruct the model to output a short answer without any intermediate reasoning or evidence selection. SimpleDeepSearcher Sun et al. (2025b) constructs a high-quality dataset containing intermediate reasoning and retrieval steps, then fine-tunes the model with question-to-(reasoning,answer) pairs. Its input is also only the question, but its output format follows R1-Searcher with four components:
  • <thinking> … </thinking>,
  • <|begin_search_query|> … </|end_search_query|>,
  • <|begin_search_result|> … </|end_search_result|>,
[…]
(4) RL methods: Search-R1 Jin et al. (2025a) uses only F1 as reward. R1-Searcher Song et al. (2025) uses answer reward plus a format reward. Research Chen et al. (2025) also incorporates format reward. To ensure fairness, we re-trained these RL-based baselines on our training set using the authors’ released code and hyperparameter settings, without using their checkpoints. LeTS Zhang et al. (2025a) combines step-level rewards with answer rewards and introduces rollout-level redundancy penalties and group-level knowledge-matching rewards. Since LeTS has no public code, we report the results from the original paper.
All baselines are evaluated under a unified protocol: each method first produces its predicted answers, which are then scored using the same standardized evaluation script.
Figure 3: Effect of different hyperparameters.
Appendix D Detailed Reinforce++
Reinforce++ Hu et al. (2025a) is an efficient RLHF algorithm without a critic network, designed to address overfitting in advantage estimation and reward hacking in REINFORCE-based methods. Its core idea is to use the global batch mean reward as the baseline, rather than constructing a separate baseline for each prompt as in RLOO or GRPO. This avoids prompt-specific bias and improves stability and generalization. We adopt the same reward computation strategy as in our main experiments to ensure consistency across training and evaluation.
Appendix E Visualization of Annealing
Figure 4: Visualization of different reward weight scheduling strategies curves.
To better illustrate the dynamic weight schedule defined in the main content, Figure 4 plots the four annealing strategies used in Section 7.2. Our strategy keeps the weight stable in the early and mid training phases, and then drops sharply in the late phase, aligning with our goal of letting the model focus on answer accuracy in the later stages.
Method Hotpotqa 2wikimultihopqa Musique Bamboogle
EM F1 CEM EM F1 CEM EM F1 ACC CEM F1 CEM
Naive GRPO 36.4 48.5 41.6 46.4 53.1 54.0 16.2 26.0 19.6 36.0 49.2 40.8
TIRESRAG-R1-Instruct 41.0 54.2 46.0 52.8 59.6 60.8 19.4 30.0 23.2 44.0 54.7 47.2
w/o Filter 18.8 24.6 26.8 21.4 26.5 28.8 6.0 11.4 9.4 19.2 28.8 25.6
w/o Difficulty 38.2 50.4 43.6 49.2 54.0 55.4 17.0 27.0 21.4 35.2 49.3 38.4
w/o Penalty 38.0 49.9 44.2 44.2 50.9 52.0 15.6 24.9 19.0 39.2 50.6 43.2
w/o Reflect 37.8 47.9 41.2 44.6 52.0 53.8 10.8 21.4 15.2 32.8 45.3 37.6
w/o Sufficient 37.4 48.5 44.6 41.4 46.3 46.4 14.0 22.3 16.4 32.6 43.4 35.8
w/o Thinking 39.8 51.8 47.6 44.8 51.3 53.8 14.8 23.9 19.2 37.6 46.4 39.2
Table 7: Ablation Study.
Appendix F Ablation Study
To verify the effectiveness of each module in TIRESRAG-R1, we conduct systematic ablation experiments on the Qwen2.5-3B-Instruct model. Table 7 shows the obtained results. We observe that: (1) The filter module is crucial for model stability. As shown in Table 7, removing it significantly degrades performance: compared with standard GRPO, the average F1 score drops by 21.2%. This is because, in the later training stages, the model collapses. We show the training curves in Section 7.3. (2) The difficulty-aware weighting and penalty mechanisms are crucial for effectively integrating our diverse reward signals. While ablations without them still outperform naive GRPO, the gains are limited, showing average performance drops of 4.48 and 5.58 points compared to our full method. (3) Each reward component plays a critical role in overall performance. Removing any single reward leads to noticeable degradation—and in all cases, performance drops below that of naive GRPO. Notably, removing the sufficient reward results in the largest decline, indicating that the model may engage in reward hacking and neglect crucial external documents. In contrast, removing the thinking reward has the smallest impact, with an average drop of 6.3% compared to our full method. This is likely because the answer, sufficient, and reflect rewards already provide partial supervision for generating high-quality reasoning chain.
Appendix G Analysis
G.1 Impact of Hyperparameters.
We explore three key hyperparameters: the number of retrieved documents, the penalty weight, and the reward mixture weights. Figure 3 present the obtained results.
  • Impact of number of retrieved documents. As the number of retrieved documents increases from 3 to 5 (ours), the model performance improves, with EM increasing by 3.95 points on average. When increasing from 5 (ours) to 7, the average EM drops by 3.6 points, suggesting that retrieving too many documents introduces noise, hurting reasoning quality. There is thus a trade-off between providing more information and avoiding noise.
  • Impact of penalty weight. In Eq. 7, $\lambda_{p}$ controls the magnitude of the consistency penalty in the advantage calculation. Setting $\lambda_{p}=0.5$ or $\lambda_{p}=1$, we observe that as $\lambda_{p}$ increases, performance decreases, with the largest drop when $\lambda_{p}=1$. This suggests over-emphasizing consistency can suppress beneficial reward signals and hurt final performance.
  • Impact of reward mixture weights. In Eq. 5, thinking, sufficient, and reflect rewards are combined with different weights. We try settings of (0.3,0.3,0.6), (0.3,0.6,0.3), (0.3,0.3,0.3), and our (0.6,0.3,0.3). Results show that giving the highest weight to thinking reward (ours) yields the best performance, likely because thinking directly measures the quality of the reasoning chain, which more strongly impacts final answers.
Figure 5: Comparison of different reward weight scheduling strategies.
G.2 Impact of Annealing Strategy.
As described in Section 4.3, we adopt a decaying schedule for mixed reward weights over training steps. To analyze its effectiveness, we compare several variants (fixed weights, linear decay, fast decay). Weight curves are shown in Fig. 4. Results in Fig. 5 show that linear annealing performs worst, with an average EM drop of 2.35 points across datasets, suggesting auxiliary signals decay too early. In contrast, our proposed schedule achieves the best or second-best results on all datasets, especially on 2WikiMultiHopQA and Bamboogle. Fixed scheduling is relatively stable but suboptimal. Pre-low annealing performs slightly better on HotpotQA but worse on others.
Figure 6: Training dynamics of answer rewards over steps.
Figure 7: Training dynamics of thinking rewards over steps.
Method Nq Popqa Triviaqa
EM F1 CEM EM F1 CEM EM F1 CEM
Direct Generation 6.8 10.8 9.5 8.6 11.8 9.4 7.5 19.0 9.5
COT 10.5 17.5 15.0 9.7 13.8 10.7 7.8 20.6 10.1
Naive RAG 23.4 32.9 31.9 29.4 37.0 34.5 12.9 29.9 16.5
Sure 25.5 34.2 27.9 30.4 35.7 31.1 13.4 29.7 16.0
IRCOT 15.4 25.1 33.2 25.1 31.5 35.8 10.1 25.0 17.6
Self-ask 17.2 27.1 41.4 24.8 31.7 41.5 9.5 24.1 20.4
SFT 6.3 12.9 10.1 7.6 11.8 8.4 5.2 14.9 7.0
SimpleDeepSearcher 33.4 44.0 44.1 38.9 44.3 44.1 59.6 67.2 66.5
ReSearch-Instruct 35.8 46.2 44.5 41.8 47.4 46.3 58.4 66.1 64.1
Search-R1-Instruct 34.2 44.0 44.2 37.9 43.5 44.1 54.5 62.2 62.3
R1-search-Instruct 35.2 46.4 44.8 40.3 46.1 45.6 57.3 65.6 64.0
TIRESRAG-R1-Instruct 38.0 49.1 47.9 43.0 48.8 47.9 60.0 68.2 66.9
Table 8: Generalization results on single-hop benchmarks (NQ, PopQA, TriviaQA).
Figure 8: Comparison of different reward backbone models.
G.3 Learning Curve Analysis.
Figures 6 and 7 show training dynamics. For naive GRPO, because it does not filter out “all-correct” or “all-wrong” queries, its answer reward is similar to ours early on but then drops sharply, showing clear collapse, before slowly recovering but remaining below initial levels. Our method maintains stability throughout training, with answer reward staying high. For the baseline without thinking reward, answer reward remains similar to ours, but thinking reward fluctuates and declines, indicating the model fails to learn a stable reasoning process. This further shows that process-level rewards help both reasoning quality and training stability.
G.4 Influence of Reward Backbone Model.
The reward model plays a critical role. Fig. 8 compares using Qwen2.5-3B and Qwen3-3B for sufficient and thinking scoring. We observe consistent improvements with Qwen3-3B: on Musique, EM improves from 16.6 to 17.6; on Bamboogle, EM improves from 41.6 to 42.8. A similar trend is seen comparing Qwen2.5-3B and Qwen2.5-7B: on HotpotQA, EM improves from 39.4 to 40.4 and F1 from 52.7 to 54.4. These results show that stronger reward models provide better supervision and thus improve final performance.
Appendix H Generalization on Single-Hop Benchmarks
We present in Table 8 the complete experimental results with more baselines and additional metrics. As shown, our method outperforms all baselines across all metrics.
Appendix I Case Study
Figures 14, 15, and 16 present three examples illustrating the effectiveness of TIRESRAG-R1. In the first example, the TIRESRAG-R1-trained Qwen2.5-3B follows a correct reasoning process and gives the correct answer, while the naive GRPO model, despite having sufficient information, produces an incorrect reasoning chain and thus a wrong answer. In the second example, the TIRESRAG-R1-trained model successfully retrieves enough information before answering, whereas the naive GRPO model guesses prematurely without sufficient evidence and produces a wrong answer. In the third example, the naive GRPO model has sufficient information and correct reasoning but still outputs the wrong answer, while our model successfully produces the correct answer. These examples show that TIRESRAG-R1 guides models toward better reasoning chains and ensures stronger consistency between reasoning and answers.
Appendix J Prompt Templates
Figures 9 to 13 present all prompt templates mentioned in this paper.
  • Figure 9: Prompt used for training instruction-based models, setting the system’s reasoning–retrieval–reflection strategy and output format.
  • Figure 10: Prompt used for training base models, converting the above instruction-based prompt into a conversation format.
  • Figure 11: Prompt for scoring the sufficiency of reasoning trajectories. We follow the sufficient-context method: the model is prompted to list sub-questions that solve the main question, and only if all sub-questions can be answered from the given references is the trajectory judged sufficient. A demonstration example is provided.
  • Figure 12: Prompt for scoring reasoning quality. The four criteria described in the main text (logical soundness, alignment, error awareness, conciseness) are included. The model is prompted to output only a number between 0 and 1 to avoid format errors. In experiments, the reward model followed the prompt strictly and output valid scores.
  • Figure 13: Prompt used for LLM-as-Judge evaluation. We embed both the predicted answer and the gold answer into the prompt and feed it to GPT-4o. If GPT-4o judges that the predicted and gold answers are semantically equivalent, it returns “yes,” otherwise “no.”
Appendix K Algorithm
The pseudo code for TIRESRAG-R1 is shown in Algorithm 1.
Algorithm 1 TIRESRAG-R1 Training with GRPO and Multi-Dimensional Rewards
0: Policy model $\pi_{\theta}$, reference model $\pi_{\theta_{\text{old}}}$, dataset $\mathcal{D}$, retrieval model $\pi_{\text{ret}}$, hyperparameter weights $w_{\text{think}},w_{\text{suff}},w_{\text{reflect}}$, KL penalty $\beta$, iterations $T$, rollouts per query $G$, buffer $B$, dynamic weight $d_{w}$
0: Updated policy model $\pi_{\theta}$
1:for $t=1,\dots,T$ do
2:  Compute dynamic weight:
$a_{t}\leftarrow\frac{1}{1+\exp\!\left(\frac{T-0.9t}{10}\right)}$
3:  Sample $Q\subset\mathcal{D}$
4:  for $q\in Q$ do
5:   Generate rollouts $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid q)$
6:   for $i=1,\dots,G$ do
7:    Extract reasoning trajectory $RD_{i}$ and prediction $a_{i}$ from $y_{i}$
8:    Compute answer reward:
$R_{i}^{A}\leftarrow\mathrm{F1}(a_{i},a^{*})$
where $a^{*}$ is the gold answer for $q$
9:    Compute sufficient reward:
$\mathrm{R^{S}_{i}}(q,RD_{i},a^{*})=\begin{cases}1,&RD_{i}\text{ contains sufficient info to derive }a^{*},\\ 0,&\text{otherwise}\end{cases}i$
10:    Compute think reward:
$R_{i}^{T}\leftarrow\mathrm{Think}(RD_{i}),\quad\mathrm{Think}(RD_{i})\in[0,1]$
scored on logic, alignment, error-awareness, and conciseness
11:    Compute reflection reward:
$\mathrm{R^{T}_{i}}(a_{i})=\begin{cases}+1,&\text{if }CEM(a_{1},a^{*})=0\text{ and }CEM(a_{2},a^{*})=1,\\ -1,&\text{if }CEM(a_{1},a^{*})=1\text{ and }CEM(a_{2},a^{*})=0,\\ 0,&\text{otherwise}\end{cases}$
12:    Combine rewards:
$R^{sum}_{i}=a_{t}\cdot\bigl{(}w_{\text{think}}R_{i}^{T}+w_{\text{suff}}R_{i}^{S}+w_{\text{reflect}}R_{i}^{R}\bigr{)}+R_{i}^{A}$
13:   end for
14:   if $0.1<\forall\ r_{i}^{(a)}<0.9$ for $i\in\{1,\dots,G\}$ then
15:    Compute advantage by normalizing batch reward:
$A_{i}\leftarrow\frac{R_{i}^{\mathrm{sum}}-\frac{1}{G}\sum_{j=1}^{G}R_{j}^{\mathrm{sum}}}{\sqrt{\frac{1}{G}\sum_{j=1}^{G}\left(R_{j}^{\mathrm{sum}}-\frac{1}{G}\sum_{k=1}^{G}R_{k}^{\mathrm{sum}}\right)^{2}}}$
16:    Compute consistency penalty $A_{i}^{P}$ by Eq. 7
17:    Apply difficulty-aware weighting:
[…]
18:    Add sample to buffer:
$B\leftarrow B\cup\{(q,y_{i},A_{i}^{\prime})\}_{i=1}^{G}$
19:   else
20:    continue
21:   end if
22:   Update $\pi_{\theta}$ on buffer $B$
23:  end for
24:end for
25:return $\pi_{\theta}$
Prompt Template for RAG-reasoning Model Generation for Instruction-based Model You are a helpful assistant. Answer the given question. You must reason **clearly and completely** inside <think> and </think> before providing any final answer. Always identify and verify all key entities (e.g., person names, locations, dates, awards) mentioned in the question. If you are uncertain about an entity or fact, or if the question requires external knowledge, you may use <search>your query</search>, and the top search results will be returned between <information> and </information>. Carefully read and reflect on each newly retrieved piece of information. You can search as many times as you want. When reasoning, you must ensure your reasoning path aligns strictly with the evidence. After reasoning, before providing your final answer, rethink it to make sure the answer is exactly correct for the original question. Use the most accurate span from the evidence when possible. Only after satisfying all the above, give the final answer inside <answer> and </answer>. For example, <answer>Beijing</answer>. After outputting the final answer in <answer></answer>, you have one chance to reflect on the answer. If you choose not to reflect, nothing further needs to be done. Otherwise, you can then re-examine your thinking process, the information obtained, and even search for more information to verify your previous answer or correct the previous answer. Remember, after reflection ends, you should output the answer in <answer></answer>.
Figure 9: Prompt template for instruction model training.
Prompt Template for RAG-reasoning Model Generation for Base Model Answer the given question. You must reason **clearly and completely** inside <think> and </think> before providing any final answer. Always identify and verify all key entities (e.g., person names, locations, dates, awards) mentioned in the question. If you are uncertain about an entity or fact, or if the question requires external knowledge, you may use <search>your query</search>, and the top search results will be returned between <information> and </information>. Carefully read and reflect on each newly retrieved piece of information. You can search as many times as you want. When reasoning, you must ensure your reasoning path aligns strictly with the evidence. After reasoning, before providing your final answer, rethink it to make sure the answer is exactly correct for the original question. Use the most accurate span from the evidence when possible. Only after satisfying all the above, give the final answer inside <answer> </answer>. For example, <answer> Beijing </answer>. After outputting the final answer in <answer> </answer>, you have one chance to reflect on the answer. If you choose not to reflect, nothing further needs to be done. Otherwise, you can then re-examine your thinking process, the information obtained, and even search for more information to verify your previous answer or correct the previous answer. Remember, after reflection ends, you should output the answer in <answer> </answer>.
Figure 10: Prompt template for base model training.
Prompt for Sufficient Reward Evaluation You are an expert LLM evaluator that excels at evaluating a QUESTION, ANSWER and REFERENCES. Consider the following criteria:
Sufficient Context To The Given Answer: 1 IF the CONTEXT is sufficient to infer the ANSWER to the question and 0 IF the CONTEXT cannot be used to infer the ANSWER to the question. Make the sufficiency judgment based solely on the context, without relying on your memory to determine whether the question can be answered from the context.
First, output a list of step-by-step questions that would be used to arrive at a label for the criteria. Make sure to include questions about assumptions implicit in the QUESTION. Include questions about any mathematical calculations or arithmetic that would be required.
Next, answer each of the questions. Please note that you may answer these questions only on the basis of the given context; do not use your own outside knowledge. Make sure to work step by step through any required mathematical calculations or arithmetic. Finally, use these answers to evaluate the criteria.
EVALUATION (JSON)
EXAMPLE:
### QUESTION
In which year did the publisher of Roald Dahl’s Guide to Railway Safety cease to exist?
### ANSWER
2001
### References
Roald Dahl’s Guide to Railway Safety was published in 1991 by the British Railways Board. The British Railways Board had asked Roald Dahl to write the text of the booklet, and Quentin Blake to illustrate it, to help young people enjoy using the railways safely. The British Railways Board (BRB) was a nationalised industry in the United Kingdom that operated from 1963 to 2001. Until 1997 it was responsible for most railway services in Great Britain, trading under the brand name British Railways and, from 1965, British Rail. It did not operate railways in Northern Ireland, where railways were the responsibility of the Government of Northern Ireland.
### EXPLANATION
The context mentions that Roald Dahl’s Guide to Railway Safety was published by the British Railways Board. It also states that the British Railways Board operated from 1963 to 2001, meaning the year it ceased to exist was 2001. Therefore, the context does provide a precise answer to the question.
### JSON
{{"Sufficient Context To The Given Answer": 1}}
Remember the instructions: You are an expert LLM evaluator that excels at evaluating a QUESTION, ANSWER and REFERENCES. Consider the following criteria:
Sufficient Context: 1 IF the CONTEXT is sufficient to infer the ANSWER to the question and 0 IF the CONTEXT cannot be used to infer the ANSWER to the question. Make the sufficiency judgment based solely on the context, without relying on your memory to determine whether the question can be answered from the context.
First, output a list of step-by-step questions that would be used to arrive at a label for the criteria. Make sure to include questions about assumptions implicit in the QUESTION. Include questions about any mathematical calculations or arithmetic that would be required.
Next, answer each of the questions. Please note that you may answer these questions only on the basis of the given context; do not use your own outside knowledge. Make sure to work step by step through any required mathematical calculations or arithmetic.
Finally, use these answers to evaluate the criteria. Output the ### EXPLANATION (Text). Then, use the EXPLANATION to output the ### EVALUATION (JSON)
Figure 11: Prompt used to evaluate context sufficiency.
Prompt for Thinking Reward Evaluation You are an expert reasoning evaluator for Retrieval-Augmented Generation (RAG) tasks.
Your goal is to judge the reasoning quality of the model’s thinking process based on the retrieved context and question.
You will assign a reward score between 0 and 1. This score reflects only the quality of the reasoning process, not whether the final answer is correct.
Evaluation Criteria:
1. Logical Soundness – Is the reasoning coherent and structured?
2. Contextual Alignment – Does it use retrieved evidence correctly?
3. Error Awareness – Does it avoid unsupported assumptions?
4. Clarity and Precision – Is it concise, relevant, and non-redundant?
Scoring:
0.0: Completely flawed reasoning
1.0: Perfect reasoning
Intermediate (e.g., 0.3, 0.7) are allowed.
Important:
- Judge only the thinking process, not the answer.
- Reward accurate, grounded, and structured reasoning.
Your Output:
A single float-type score from {{0.0, 0.1, 0.2, ..., 1.0}}.
No explanation. Only the score.
Figure 12: Prompt used to evaluate thinking quality
Prompt for LLM-As-Judge Given a Question and its Golden Answer, verify whether the Predicted Answer is correct.
The prediction is correct if it fully aligns with the meaning and key information of the Golden Answer.
Respond with True if the prediction is correct and False otherwise.
Question: {question}
Golden Answer: {reference}
Predicted Answer: {prediction}
Figure 13: Prompt used for LLM-as-Judge evaluation.
Question: Who is Gaius Rubellius Blandus’s father-in-law? Golden Answer: Drusus Julius Caesar Naive GRPO: <think> To find out who Gaius Rubellius Blandus’s father-in-law is, I need to first determine who his wife was. I’ll search for information about Gaius Rubellius Blandus to find details about his family, particularly his wife. </think>
<search>
Gaius Rubellius Blandus family </search>
<information>
… (information retrieved about Julia Livia being granddaughter of Tiberius, but not focusing on her father) … </information>
<think>
From the information retrieved, we know that Gaius Rubellius Blandus married Julia in AD 33. Now, I need to find out who Julia’s father was to determine her father-in-law. </think>
<search>
Julia Livia Gaius Rubellius Blandus wife </search>
<information>
… (mentions Julia Livia is granddaughter of Tiberius) … </information>
<think>
From the information retrieved, we know that Julia Livia was the granddaughter of Tiberius. Therefore, Gaius Rubellius Blandus’s father-in-law is Tiberius. </think>
<answer>
Tiberius </answer>
TIRESRAG-R1: <think> First, I need to identify who Gaius Rubellius Blandus’s wife was. Then I need to find out who her father was. </think>
<search>
who was Gaius Rubellius Blandus’s wife </search>
<information>
Gaius Rubellius Blandus married Julia Livia, granddaughter of Emperor Tiberius, in AD 33. … </information>
<think>
Now that I know his wife is Julia Livia, I need to find out who Julia Livia’s father was. </think>
<search>
who was Julia Livia’s father </search>
<information>
Julia Livia was the daughter of Drusus Julius Caesar and Livilla. … </information>
<think>
Therefore, Gaius Rubellius Blandus’s father-in-law is Drusus Julius Caesar. </think>
<answer>
Drusus Julius Caesar </answer>
Figure 14: Case study where the naive GRPO model, despite having sufficient information, mistakenly infers the grandfather Tiberius as the father-in-law, while TIRESRAG-R1 correctly reasons through the steps and outputs Drusus Julius Caesar.
Question: Where was the place of burial of John De Ferrers, 1st Baron Ferrers of Chartley’s father? Golden Answer: Staffordshire Naive GRPO: <think> To find out where John de Ferrers, 1st Baron Ferrers of Chartley’s father was buried, I need to first determine who his father was. I’ll search for information about John de Ferrers, 1st Baron Ferrers of Chartley to find his father’s name. </think>
<search>
John de Ferrers, 1st Baron Ferrers of Chartley father </search>
<information>
… (information retrieved about Roger Ferrers of Fiddington Gloucestershire, unrelated to the correct father) … </information>
<think>
From the information retrieved, we know that John de Ferrers, 1st Baron Ferrers of Chartley’s father was Roger Ferrers of Fiddington Gloucestershire. Now, I need to find out where Roger Ferrers of Fiddington Gloucestershire was buried. </think>
<search>
Roger Ferrers of Fiddington Gloucestershire burial place </search>
<information>
… (mentions Fiddington and other unrelated burial info) … </information>
<think>
Based on the information retrieved, Roger Ferrers of Fiddington Gloucestershire was buried at Fiddington. Therefore, the place of burial of John de Ferrers, 1st Baron Ferrers of Chartley’s father was Fiddington, Gloucestershire. </think>
<answer>
Fiddington, Gloucestershire </answer>
TERRISRAG-R1: <think> First, we need to determine who John de Ferrers, 1st Baron Ferrers of Chartley’s father was. Then we need to find out where his father was buried. </think>
<search>
John de Ferrers, 1st Baron Ferrers of Chartley father burial place </search>
<information>
John de Ferrers, 1st Baron Ferrers of Chartley (1271–1324) was the son of Robert de Ferrers, 6th Earl of Derby and Alianore de Bohun. … </information>
<think>
From the search results, I can confirm that John de Ferrers, 1st Baron Ferrers of Chartley’s father was Robert de Ferrers, 6th Earl of Derby. Now I need to find where he was buried. </think>
<search>
Robert de Ferrers, 6th Earl of Derby burial place </search>
<information>
Robert de Ferrers, 6th Earl of Derby, was buried at St Thomas’s Priory at Stafford, in Staffordshire. …</information>
<think>
Therefore, the place of burial of John de Ferrers, 1st Baron Ferrers of Chartley’s father is Stafford, in Staffordshire. </think>
<answer>
Stafford, in Staffordshire </answer>
Figure 15: Case study where the naive GRPO model, despite retrieving sufficient information, follows an incorrect reasoning chain by focusing on an unrelated person (Roger Ferrers of Fiddington) and outputs the wrong burial place. In contrast, TIRESRAG-R1 reasons through the correct lineage and finds the true location in Staffordshire.
Question: Who was born first, Dennis E. Nolan or Humberto Anguiano? Golden Answer: Dennis E. Nolan Naive GRPO: <think> To determine who was born first between Dennis E. Nolan and Humberto Anguiano, I need to find out their birth dates. I’ll first search for Dennis E. Nolan’s birth date. </think>
<search>
Dennis E. Nolan birth date </search>
<information>
… Dennis E. Nolan (1872–1956), United States Army general. … </information>
<think>
From the search results, I found that Dennis E. Nolan was a United States Army general. Now, I’ll search for Humberto Anguiano’s birth date. </think>
<search>
Humberto Anguiano birth date </search>
<information>
… Humberto Anguiano (born 2 November 1910) was a Mexican modern pentathlete. …</information>
<think>
Based on the information retrieved, Dennis E. Nolan was born in 1872 while Humberto Anguiano was born in 1910. Therefore, Dennis E. Nolan was born first. </think>
<answer>
Humberto Anguiano </answer>
TERRISRAG-R1: <think> Let’s determine who was born first, Dennis E. Nolan or Humberto Anguiano, by finding their respective birth years. </think>
<search>
Dennis E. Nolan birth date </search>
<information>
… Dennis E. Nolan (1872–1956), United States Army general. … </information>
<think>
From the search results, I can confirm that Dennis E. Nolan is a United States Army general born in 1872. </think>
<search>
Humberto Anguiano birth date </search>
<information>
… Humberto Anguiano (born 2 November 1910) was a Mexican modern pentathlete. … </information>
<think>
Since Dennis E. Nolan was born in 1872 and Humberto Anguiano was born in 1910, Dennis E. Nolan was born first. </think>
<answer>
Dennis E. Nolan </answer>
Figure 16: Case study where the naive GRPO model, despite retrieving the correct birth years (1872 for Dennis E. Nolan and 1910 for Humberto Anguiano), produces an inconsistent reasoning chain and wrongly outputs Humberto Anguiano as born first. In contrast, TIRESRAG-R1 follows a correct reasoning process and gives the right answer, Dennis E. Nolan.